frostyhealer:

What is the solution of the system defined by y = –x 5 and 5x 2y = 14?

K1NGofPadlet:

To solve the system defined by y = –x + 5 and 5x + 2y = 14, we can substitute the value of y from the first equation into the second equation. This gives us: 5x + 2(-x+5) = 14 Simplifying this equation, we get: 5x - 2x + 10 = 14 Combining like terms, we get: 3x + 10 = 14 Subtracting 10 from both sides, we get: 3x = 4 Dividing both sides by 3, we get: x = 4/3 Now that we know the value of x, we can use either of the original equations to find the value of y. Using y = –x + 5, we get: y = -4/3 + 5 Simplifying this expression, we get: y = 11/3 Therefore, the solution to the system defined by y = –x + 5 and 5x + 2y = 14 is (4/3, 11/3).

frostyhealer:

thank you very much!

K1NGofPadlet:

@frostyhealer wrote:
thank you very much!
Of course 😉